The basement for a house is being excavated and the rectangular hole measured 11m by 14,5m by 4m. The soil is being hauled away in a dump truck which has a an open- box bed of 4m by 3m by 200cm. The truck can't be overloaded at anytim. How many trips will the dump truck have to make to haul away all the soil?


1
Expert's answer
2022-02-21T17:52:22-0500

"\text{Number of trips required} =\dfrac{\text{Volume of the soil excavated}}{\text{Volume of the dump truck}}"

Given, the measurements of the rectangular base for the house being excavated and the measurements of the dump truck.

Here, 200cm = 2m.

Therefore,


"\text{Number of trips required} =\dfrac{11\times14.5\times4}{4\times3\times2}= \dfrac{638}{24}=26.58\simeq 27"

Hence we require 27 trips to haul away the soil.